Find the flux through a spherical Gaussian surface of radius a = 1 m surrounding a charge of 8.85 pC.


1
Expert's answer
2020-12-25T14:07:06-0500

"\Phi=\frac{1}{\epsilon_0}\sum q_i=\frac{1}{\epsilon_0}q=\frac{8.85\cdot10^{-12}}{8.85\cdot10^{-12}}=1(V\cdot m)"